# Distilling Diffusion Models into Conditional GANs

### v1

@@ 1 Introduction
One simple solution is to just train a one-step model from scratch. For example, we can train a GAN [16], which stands out as a leading one-step model for simple domains [36]. Unfortunately, training text-to-image GANs on large-scale, diverse datasets is still challenging, despite recent advances [33, 85]. The challenge lies in GANs needing to tackle two difficult tasks all at once without any supervision: (1) finding correspondence between noises and natural images, and (2) effectively optimizing a generator model to perform the mapping from noises to images. This “unpaired” learning is often considered more ill-posed, as mentioned in CycleGAN [111], compared to paired learning, where conditional GANs [29] can learn to map the input to output, given ground truth correspondences.
@@
First, in regression tasks, using perceptual losses (such as LPIPS [106]) better preserves perceptually important details over point-based losses (such as L2), However, perceptual losses are fundamentally incompatible with Latent Diffusion Models [77], as they require an expensive decoding from latent to pixel space. To overcome this, we propose LatentLPIPS, showing that perceptual losses can directly work in latent space. This enables a fourfold increase in batch size, compared to computing LPIPS in pixel space. Unfortunately, we observe that the latent-based perceptual loss has more blind spots than its pixel counterparts. While previous work has found that ensembling is helpful for pixel-based LPIPS [38], we find that it is critical for the latent-based version. Working in latent space with our Ensembled-LatentLPIPS, we demonstrate strong performance with just a regression loss, comparable to guided progressive distillation [60]. Additionally, we employ a discriminator in the training to further improve performance. We develop a multi-scale conditional diffusion discriminator, leveraging the pre-trained weights and using our new single-sample R1 loss and mix-and-match augmentation. We name our distillation model Diffusion2GAN.
@@
Figure 1: Visual comparison to SDXL teacher [79] (classifier-free guidance scale [25] of 7) and selected distillation student models, including SDXL-Turbo [86], SDXL-Lightning [48], and our SDXL-Diffusion2GAN. All images in a given row were generated using the same noise input, except for SDXL-Turbo, which requires a distinct noise size of $4\times 64\times 64$. Compared to other distillation models, our SDXL-Diffusion2GAN more closely adheres to the original ODE trajectory.
@@ 2 Related Work
Diffusion distillation. Accelerating the sampling speed of diffusion models is crucial for enhancing practical applications, as well as reducing energy costs for inference. Several works have proposed reducing the number of sampling steps using fast ODE solvers [54, 55, 34] or reducing the computational time per step [45, 46, 8]. Another effective method for acceleration is knowledge distillation [22, 56, 100, 52, 83, 60, 3, 17, 110, 53, 93]. In this approach, multiple steps of a teacher diffusion model is distilled into a fewer-step student model. Luhman et al. [56] propose $L_{p}$ regression training between student’s output from a Gaussian noise $\mathbf{x}_{T}$ and its corresponding ODE solution $\mathbf{x}_{0}$. Despite its simplicity, such direct regression produces blurry outputs and does not match the image synthesis capabilities exhibited by other generative models. To enhance image quality, InstaFlow [53] straightens high-curvature ODE trajectory via ReFlow [52] and distills the linearized ODE trajectory to the student model. Consistency Distillation (CD) [93, 57] trains a student model to predict a consistent output for any noisy sample $\mathbf{x}_{t+1}$ and its single-step denoising $\mathbf{x}_{t}$. Recently, several studies have proposed using a GAN discriminator to enhance distillation performance. For example, CTM [39] and SDXL-Turbo [86] utilize an improved StyleGAN [87, 85] discriminator to train a one-step image generator. In addition, UFOGen [102], SDXL-Lightning [48], and ADD [84] adopt a pre-trained diffusion model as a strong discriminator, demonstrating their capabilities in one-step text-to-image synthesis. Although these works are concurrent, we will compare our method with SDXL-Turbo and SDXL-Lightning, both of which have demonstrated state-of-the-art performance in diffusion distillation for one-step image synthesis.
Conditional Generative Adversarial Networks [62, 29] have been a commonly-used framework for conditional image synthesis. The condition could be an image [29, 111, 9, 51, 66, 76], class-label [65, 63, 5, 37, 31], and text [75, 104, 101, 85, 33]. In particular, cGANs have shown impressive performance when helped by a regression loss to stabilize training, as in image translation [29, 97, 111, 67, 107, 68]. Likewise, we approach diffusion model distillation by employing the image-conditional GAN, along with a perceptual regression loss [106]. Early works combine GANs with the forward diffusion process [100, 99], but do not aim at distilling a pre-trained diffusion model into a GAN.
@@ 3.3 Conditional Diffusion Discriminator
Initialization from a pre-trained diffusion model. We demonstrate that initializing the discriminator weights with a pre-trained diffusion model is effective for diffusion distillation. Compared to the implementation of GigaGAN discriminator [33], using a pre-trained Stable Diffusion 1.5 U-Net [79] and finetuning the model as the discriminator in the latent space results in superior FID in Table 2. The adversarial loss is computed independently at each location of the U-Net discriminator output. Note that the original U-Net architecture conditions on text but not on the input noise map ${\mathbf{z}}$. We further modify the discriminator architecture to support ${\mathbf{z}}$ conditioning, simply by adding the input with ${\mathbf{z}}$ processed through a single convolution layer with zero initialization in the channel dimension. Note that the text conditioning for the diffusion discriminator is naturally carried out by the built-in cross-attention layers in the SD U-Net. We observe moderate improvement across all metrics.
@@ 4 Experiments
Here, we first study the effectiveness of our algorithmic designs with a systematic ablation study in Section 4.1. We then compare our method with leading one-step generators using a standard benchmark regarding image quality, text alignment, and inference speed in Section 4.2. We further provide visual comparisons (Section 4.3) and report the training speed (Section 4.4).
Training details. We distill Stable Diffusion 1.5 into our one-step generator and train the model on two ODE datasets with different classifier-free guidance (CFG), namely, the SD-CFG-3 dataset with 3 million noise-latent pairs and the SD-CFG-8 dataset with 12 million pairs. We use the prompts from the LAION-aesthetic 6.25 and 6.0 datasets for the SD-CFG-3 and SD-CFG-8 datasets, respectively, and simulate the ODE using 50 steps of DDIM sampler [91]. To demonstrate the effectiveness of Diffusion2GAN for a larger text-to-image model, we distill SDXL-Base-1.0 [71] into Diffusion2GAN using 8 million noise-latent pairs named SDXL-CFG-7 dataset. These pairs were generated by SDXL-Base-1.0 using prompts from the LAION 6.0 dataset. We simulate the ODE of SDXL-Base-1.0 using 50 steps of DDIM sampler. For further details on hyperparameters, please refer to the Appendix 0.A. Notably, we did not use any real images from the LAION dataset.
@@ 4.2 Comparison with Distilled Diffusion Models
Distilling Stable Diffusion 1.5. We compare Diffusion2GAN with cutting-edge diffusion distillation models on both COCO2014 and COCO2017 benchmarks in Tables 3 and 4, respectively. While InstaFlow-0.9B can attain an FID of 13.10 on COCO2014 and 23.4 on COCO2017, Diffusion2GAN can more efficiently deal with the ODE distillation problem, achieving an FID of 9.29 and 19.5, respectively. Similar to our method, UFOGen [102], MD-UFOGen [109], DMD [103], and ADD-M [86] use extra diffusion models for adversarial training or distribution matching. Although these models achieve lower FIDs compared to InstaFlow-0.9B, Diffusion2GAN still outperforms them, as Diffusion2GAN is trained to closely follow the original trajectory of the teacher diffusion model, thus mitigating the diversity collapse issue while maintaining high visual quality. Note that ADD-M employs a ViT-g-14 text encoder for discriminator conditioning, which is also used in CLIP-5k score calculation. The high CLIP-5k score achieved by ADD-M might be partly attributed to the use of the ViT-g-14 text encoder, whereas Diffusion2GAN is trained without it.
@@
Table 4: Comparison to recent text-to-image models on COCO2017. On the SD-CFG-3, Diffusion2GAN, distilled from Stable Diffusion 1.5 [79], demonstrates better performance over UFOGen [102]. While Diffusion2GAN presents comparable FID [21] with ADD-M [86], it exhibits worse CLIP-score [20]. Note that ADD-M is trained using the ViT-g-14 text encoder, which is also used to compute the CLIP-score.
@@ 5 Discussion and Limitations
Limitations. Although our method achieves faster inference while maintaining image quality, it does have several limitations. First, our current approach simulates a fixed classifier-free guidance scale, a common technique for adjusting text adherence, but does not support varying CFG values at inference time. Exploring methods like guided distillation [60] could be a promising direction. Second, as our method distills a teacher model, the performance limit of our model is bound by the quality of the original teacher’s output. Further leveraging real text and image pairs is a potential avenue to learn a student model that outperforms the original teacher model. Last, while Diffusion2GAN alleviates the diversity drop by introducing ODE distillation loss and a conditional GAN framework, we have found that the diversity drop still occurs as we scale up the student and teacher models. We leave further investigation of this problem for future work. Acknowledgments. We would like to thank Tianwei Yin, Seungwook Kim, Sungyeon Kim for their valuable feedback and comments. Part of this work was done while Minguk Kang was an intern at Adobe Research. Minguk Kang and Suha Kwak were supported by the NRF grant and IITP grant funded by Ministry of Science and ICT, Korea (NRF-2021R1A2C3012728, AI Graduate School (POSTECH): IITP-2019-0-01906). Jaesik Park was supported by the IITP grant funded by the government of South Korea (MSIT) (AI Graduate School (SNU): 2021-0-01343 and AI Innovation Hub: 2021-0-02068). Jun-Yan Zhu was supported by the Packard Fellowship.
@@ 0.A.1 Text-to-Image Synthesis
Two-stage Diffusion2GAN training. We observe enhanced stability and increased diversity in image generation when employing a two-stage training approach for Diffusion2GAN. In the initial stage, Diffusion2GAN is exclusively trained using the ODE distillation loss. Subsequently, we fine-tune the ODE-distilled one-step generator by incorporating the ODE distillation, conditional GAN, and single-sample R1 losses. Detailed hyperparameters are provided in Table A1.

### v2

@@ 1 Introduction
One simple solution is to just train a one-step model from scratch. For example, we can train a GAN [16], a leading one-step model for simple domains [37]. Unfortunately, training text-to-image GANs on large-scale and diverse datasets is still challenging, despite recent advances [33, 86]. The challenge lies in GANs needing to tackle two difficult tasks all at once without any supervision: (1) finding correspondence between noises and natural images, and (2) effectively optimizing a generator model to perform the mapping from noises to images. This “unpaired” learning is often considered more ill-posed, as mentioned in CycleGAN [113], compared to paired learning, where conditional GANs [29] can learn to map the input to output, given ground truth correspondences.
@@
First, in regression tasks, using perceptual losses (such as LPIPS [108]) better preserves perceptually important details over point-based losses (such as L2). However, perceptual losses are fundamentally incompatible with Latent Diffusion Models [78], as they require an expensive decoding from latent to pixel space. To overcome this, we propose LatentLPIPS, showing that perceptual losses can directly work in latent space. This enables a fourfold increase in batch size, compared to computing LPIPS in pixel space. Unfortunately, we observe that the latent-based perceptual loss has more blind spots than its pixel counterparts. While previous work has found that ensembling is helpful for pixel-based LPIPS [39], we find that it is critical for the latent-based version. Working in latent space with our Ensembled-LatentLPIPS, we demonstrate strong performance with just a regression loss, comparable to guided progressive distillation [61]. Additionally, we employ a discriminator in the training to further improve performance. We develop a multi-scale conditional diffusion discriminator, leveraging the pre-trained weights and using our new single-sample R1 loss and mix-and-match augmentation. We name our distillation model Diffusion2GAN.
@@
Figure 1: Visual comparison to SDXL teacher [72] with a classifier-free guidance scale [25] of 7 and selected distillation student models, including SDXL-Turbo [87], SDXL-Lightning [49], and our SDXL-Diffusion2GAN. All images in a given row were generated using the same noise input, except for SDXL-Turbo, which requires a distinct noise size of $4\times 64\times 64$. Compared to other distillation models, our SDXL-Diffusion2GAN more closely adheres to the original ODE trajectory.
@@ 2 Related Work
Diffusion distillation. Accelerating the sampling speed of diffusion models is crucial for enhancing practical applications, as well as reducing energy costs for inference. Several works have proposed reducing the number of sampling steps using fast ODE solvers [55, 56, 34] or reducing the computational time per step [46, 47, 8]. Another effective method for acceleration is knowledge distillation [22, 57, 101, 53, 84, 61, 3, 17, 112, 54, 94]. In this approach, multiple steps of a teacher diffusion model is distilled into a fewer-step student model. Luhman et al. [57] propose $L_{p}$ regression training between student’s output from a Gaussian noise $\mathbf{x}_{T}$ and its corresponding ODE solution $\mathbf{x}_{0}$. Despite its simplicity, such direct regression produces blurry outputs and does not match the image synthesis capabilities exhibited by other generative models. To enhance image quality, InstaFlow [54] straightens high-curvature ODE trajectory via ReFlow [53] and distills the linearized ODE trajectory to the student model. Consistency Distillation (CD) [94, 58] trains a student model to predict a consistent output for any noisy sample $\mathbf{x}_{t+1}$ and its single-step denoising $\mathbf{x}_{t}$. Recently, several studies have proposed using a GAN discriminator to enhance distillation performance. For example, CTM [40] and SDXL-Turbo [87] utilize an improved StyleGAN [88, 86] discriminator to train a one-step image generator. In addition, UFOGen [103], SDXL-Lightning [49], and LADD [85] adopt a pre-trained diffusion model as a strong discriminator, demonstrating their capabilities in one-step text-to-image synthesis. Although these works are concurrent, we will compare our method with SDXL-Turbo and SDXL-Lightning, both of which have demonstrated state-of-the-art performance in diffusion distillation for one-step image synthesis.
Conditional Generative Adversarial Networks [63, 29] have been a commonly-used framework for conditional image synthesis. The condition could be an image [29, 113, 9, 52, 67, 77], class-label [66, 64, 5, 38, 31], and text [76, 106, 102, 86, 33]. In particular, cGANs have shown impressive performance when helped by a regression loss to stabilize training, as in image translation [29, 98, 113, 68, 109, 69]. Likewise, we approach diffusion model distillation by employing the image-conditional GAN, along with a perceptual regression loss [108]. Early works [101, 100] combine GANs with the forward diffusion process, but do not aim at distilling a pre-trained diffusion model into a GAN.
@@ 3.3 Conditional Diffusion Discriminator
Initialization from a pre-trained diffusion model. We demonstrate that initializing the discriminator weights with a pre-trained diffusion model is effective for diffusion distillation. Compared to the implementation of GigaGAN discriminator [33], using a pre-trained Stable Diffusion 1.5 U-Net [80] and finetuning the model as the discriminator in the latent space results in superior FID in Table 2. The adversarial loss is computed independently at each location of the U-Net discriminator output. Note that the original U-Net architecture conditions on text but not on the input noise map ${\mathbf{z}}$. We further modify the discriminator architecture to support ${\mathbf{z}}$ conditioning, simply by adding the input with ${\mathbf{z}}$ processed through a single convolution layer with zero initialization in the channel dimension. Note that the text conditioning for the diffusion discriminator is naturally carried out by the built-in cross-attention layers in the Stable Diffusion U-Net. We observe moderate improvement across all metrics.
@@ 4 Experiments
Here, we first study the effectiveness of our algorithmic designs with a systematic ablation study in Section 4.1. Next, We compare our method with leading one-step generators using a standard benchmark regarding image quality, text alignment, and inference speed in Section 4.2. We then present human preference evaluation results in Section 4.3. Additionally, we provide visual comparisons (Section 4.4) and report the training speed (Section 4.5).
Training details. We distill Stable Diffusion 1.5 into our one-step generator and train the model on two ODE datasets with different classifier-free guidance (CFG), namely, the SD-CFG-3 dataset with 3 million noise-latent pairs and the SD-CFG-8 dataset with 12 million pairs. We use the prompts from the LAION-aesthetic-6.25 and -6.0 datasets to create the SD-CFG-3 and SD-CFG-8 datasets, respectively, and simulate the ODE using 50 steps of DDIM [92]. To demonstrate the effectiveness of Diffusion2GAN for a larger text-to-image model, we distill SDXL-Base-1.0 [72] into Diffusion2GAN using 8 million noise-latent pairs named SDXL-CFG-7 dataset. These pairs were generated by SDXL-Base-1.0 using prompts from the LAION-aesthetic-6.0 dataset. We simulate the ODE of SDXL-Base-1.0 using 50 steps of DDIM. For further details on hyperparameters, please refer to the Appendix 0.A. Notably, we did not use any real images from the LAION dataset.
@@ 4.2 Comparison with Distilled Diffusion Models
Distilling Stable Diffusion 1.5. We compare Diffusion2GAN with cutting-edge diffusion distillation models on both COCO2014 and COCO2017 benchmarks in Tables 3 and 4, respectively. While InstaFlow-0.9B can attain an FID of 13.10 on COCO2014 and 23.4 on COCO2017, Diffusion2GAN can more efficiently deal with the ODE distillation problem, achieving an FID of 9.29 and 19.5, respectively. Similar to our method, UFOGen [103], MD-UFOGen [111], DMD [104], and ADD-M [87] use extra diffusion models for adversarial training or distribution matching. Although these models achieve lower FIDs compared to InstaFlow-0.9B, Diffusion2GAN still outperforms them, as Diffusion2GAN is trained to closely follow the original trajectory of the teacher diffusion model, thus mitigating the diversity collapse issue while maintaining high visual quality. Note that the concurrent work ADD-M exhibits a higher CLIP-score compared to Diffusion2GAN. We hypothesize this is because ADD-M conditions the discriminator using both image and text embeddings, as shown in Table 1(b) of the ADD-M paper [87]. While Diffusion2GAN focuses on efficiency and prefers not to produce the pixels required for obtaining image embeddings, in the SDXL distillation experiment, we show that Diffusion2GAN demonstrates better image quality and text-to-image alignment compared to ADD-XL, also known as SDXL-Turbo.
@@
Table 4: Comparison to recent text-to-image models on COCO2017. On the SD-CFG-3, Diffusion2GAN, distilled from Stable Diffusion 1.5 [80], demonstrates better performance over UFOGen [103]. While Diffusion2GAN presents slightly better FID [21] than ADD-M [87], it exhibits a lower CLIP-score [20].
@@
4.3 Human Preference Evaluation
We conduct human preference evaluations following the procedure described in the LADD paper [85], using 128 PartiPrompts [105] to assess preferences for image realism and text-to-image alignment. We compare Diffusion2GAN with its baselines and teacher model as shown in Figure 8 and ensure a fair comparison by comparing models distilled from the same teacher model.
For Stable Diffusion 1.5 distillation, Diffusion2GAN shows better human preferences for both image realism and text-to-image alignment over InstaFlow-0.9B. For SDXL-Base-1.0 distillation, SDXL-Diffusion2GAN demonstrates comparable or superior image realism and text-to-image alignment compared to SDXL-Turbo and SDXL-Lightning. Notably, one aspect (realism or text-to-image alignment) of SDXL-Diffusion2GAN tends to be preferred over SDXL-Turbo and SDXL-Lightning when the other aspect is comparable to the baselines. While our proposed one-step generators exhibit better realism and text-to-image alignment than previous one-step models, the multi-step teacher models are still preferred overall. We leave the future improvement of Diffusion2GAN for future work.
Figure 8: Human preference evaluation. We evaluate human preferences using images generated from 128 PartiPrompts [105]. Annotators evaluate both the image quality and text-to-image alignment. All models, except for Stable Diffusion 1.5 (SD 1.5) [80] and SDXL-Base-1.0 [72], generate images using a single forward evaluation. SD 1.5 and SDXL-Base-1.0 generated images using 50 steps of DDIM [92].
4.4 Visual Analysis
@@ 5 Discussion and Limitations
Limitations. Although our method achieves faster inference while maintaining image quality, it does have several limitations. First, our current approach simulates a fixed classifier-free guidance scale, a common technique for adjusting text adherence, but does not support varying CFG values at inference time. Exploring methods like guided distillation [61] could be a promising direction. Second, as our method distills a teacher model, the performance limit of our model is bound by the quality of the original teacher’s output. To enhance the quality of generated noise-image pairs, employing advanced diffusion models like EDM2 [36], which is better compatible with deterministic sampling, could be advantageous. Additionally, leveraging real text and image pairs is a potential avenue to learn a student model that outperforms the original teacher model. Third, Diffusion2GAN only supports one-step image synthesis as it was trained to translate given noise into an RGB image directly. However, extending Diffusion2GAN to multi-step generation could result in future performance improvement. Last, while Diffusion2GAN alleviates the diversity drop by introducing ODE distillation loss and a conditional GAN framework, we have found that the diversity drop still occurs as we scale up the student and teacher models. We leave further investigation of this problem for future work.
6 Societal Impact
Our work aims to develop a one-step image synthesis framework, which could significantly improve the accessibility and affordability of generative visual models. By reducing the multi-step synthesis process into a single step, our technology promises to democratize the creation of visual content, enabling a broader range of users to harness the power of generative models for creative expression and innovation. Additionally, by reducing the need for extensive computation during both training and inference stages, our framework also helps decrease electricity usage and CO2 emissions. However, as this technology becomes more accessible, it is crucial to address concerns about potential misuse, especially in areas like sexual harassment and synthetic media manipulation.
Generative visual models have the potential to facilitate the creation of highly convincing deep fake videos and enable sophisticated impersonation techniques, presenting significant challenges for the trustworthiness of online information. Moreover, they can be utilized to generate content that may incite instances of sexual harassment. While our technology boasts compelling advantages regarding efficiency, it is imperative to acknowledge and tackle the potential societal repercussions and ethical dilemmas linked with the widespread integration of generative visual models.
Acknowledgments. We would like to thank Tianwei Yin, Seungwook Kim, and Sungyeon Kim for their valuable feedback and comments. Part of this work was done while Minguk Kang was an intern at Adobe Research. Minguk Kang and Suha Kwak were supported by the NRF grant and IITP grant funded by Ministry of Science and ICT, Korea (NRF-2021R1A2C3012728, AI Graduate School (POSTECH): IITP-2019-0-01906). Jaesik Park was supported by the IITP grant funded by the government of South Korea (MSIT) (AI Graduate School (SNU): 2021-0-01343 and AI Innovation Hub: 2021-0-02068). Jun-Yan Zhu was supported by the Packard Fellowship.
@@ 0.A.1 Text-to-Image Synthesis
Two-stage Diffusion2GAN training. We observe enhanced stability and increased diversity in image generation when employing a two-stage training approach for Diffusion2GAN. In the initial stage, Diffusion2GAN is exclusively trained using the ODE distillation loss. Subsequently, we fine-tune the ODE-distilled one-step generator by incorporating the ODE distillation, conditional GAN, and single-sample R1 losses. Experimentally, we discover that training Diffusion2GAN with a different conditional GAN loss weight typically results in stable convergence. Increasing the weight of the conditional GAN loss component enhances the fidelity of generated images but decreases image diversity. We speculate this occurs because the conditional GAN loss prioritizes realistic image synthesis over accurately learning the original ODE trajectory of the teacher model. Detailed hyperparameters are provided in Table A1.